# CTP-SCU/2025011

### v2

@@ II Entanglement entropy between two disjoint balls
Figure 3: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the solid torus CFTD. The lower half solid torus corresponds to $|\psi\rangle$. The upper half annulus corresponds to $\langle\psi|$. We calculate the entanglement entropy between two disjoint co-dimensional one balls AA and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
@@ IV Examples in CFT2 and CFT4
V Shape independence and Ryu-Takayanagi formula
@@
V.1 Shape independence
@@
V.2 Manifestation of Ryu-Takayanagi formula
Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the volume $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, up to a conformal transformation for asymmetric configurations. Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the scaled volume of the entangling region., rather than the area of the entangling surface. As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.

### v3

@@ II Entanglement entropy between two disjoint balls
Figure 3: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the solid torus CFTD. The lower half solid torus corresponds to $|\psi\rangle$. The upper half annulus corresponds to $\langle\psi|$. We calculate the entanglement entropy between two disjoint co-dimensional one balls $A$ and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
@@ IV Examples in CFT2 and CFT4
V Ryu-Takayanagi formula and Shape independence
@@
V.1 Manifestation of Ryu-Takayanagi formula
Our derivation has already manifested the RT formula. From the first line of equations (11), for a symmetric solid torus (reflection-symmetric juxtaposed configuration) as illustrated by the left panel of Figure 6, the entanglement entropy between balls $A$ and $B$ is the volume Vol(A)(A) (or equivalently Vol(B)(B)) calculated in the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ This hyperbolic volume Vol(A)(A) could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, which is precisely the entanglement wedge cross-section (EWCS), i.e RT surface, for this symmetric configuration.
As for a generic disjoint configurations in Figure 5, we simply make a conformal transformation to get the corresponding RT surface (EWCS). This actually provides a simple way to compute the minimal surface for a particular configuration. It is of interest to verify the result in a geometric picture.
It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the hyperbolic volume of the entangling region., rather than the area of the entangling surface. As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.
V.2 Shape independence